OpenStudy (anonymous):

In how many ways can a committee of 4 be chosen from a group of 9 people?

OpenStudy (campbell_st):

do you know about combinations... are are you just seeking answers so you can pass a test

OpenStudy (anonymous):

i dont know combinations

OpenStudy (er.mohd.amir):

chose 4 out of 9

OpenStudy (campbell_st):

ok... then it makes it hard to answer this question...as that's what is needed

OpenStudy (er.mohd.amir):

9C4=ans

OpenStudy (anonymous):

@Megan30400

OpenStudy (campbell_st):

so you are wanting help to understand them... or you just want answers..?

OpenStudy (anonymous):

yes i need help

OpenStudy (campbell_st):

ok... a combination is a counting technique that finds how many ways r items can be selected from a group on n items... there is a formula that uses factorial notation. the factorial notation is simply a way of writing mutiplications... and its written n! = n x (n -1) x (n -2) x .... 3 x 2 x 1 so 4! = 4 x 3 x 2 x 1 = 24 5! = 5 x 4 x 3 x 2 x 1 = 120 the formula for selecting r objects from a group of n is \[^nC_{r} = \frac{n!}{r!(n - r)!}\] so in your question n = 9 and r = 4 so you have \[^9C_{4} = \frac{9!}{4!(9 - 4)!} = \frac{9 \times 8 \times 7 \times 6 \times 5 \times 4 \times 3 \times 2 \times 1}{(4 \times 3 \times 2 \times 1)\times(5 \times 4 \times 3\times 2\times 1)}\] which can be simplified to \[^9C_{4} = \frac{9 \times8 \times 7 \times 6}{4 \times 3 \times 2 \times 1}\] which gives the answer of 126 the nice thing about recognising a combination is that you calculator has an inbuilt function that does the calculation for you.